# Optimization via Quantum Preconditioning

### v1

@@ D.2 Burer-Monteiro Solver
The BM solver eliminates the positive semi-definite constraint of the GW algorithm for the maximum-cut problem. The GW algorithm is based on semidefinite programming methods and the solver guaranteeing the highest possible approximation ratio $\alpha\simeq 0.878$ [10, 11]. Thus, the BM solver involves a nonconvex objective function which may result in multiple local nonglobal minima. In practice, this is not an issue and the BM solver is considered one of the best heuristics to date for the maximum-cut problem [12]. Its foundations are still intensively studied by the community; see, e.g., [169, 170] and references therein.
@@
We employ the MQLib implementation of the BM solver [12], written in C++ for efficiency. This implementation does not readily expose any hyperparameters for the solver and we keep the default values. That’s why, unlike for SA, the run-time is strictly proportional to the number of iterations (see Appendix H). Our only modification of the MQLib code exposes the number of iterations performed by the solver in addition to its run-time.
@@
Appendix E Terms in Quantum-Preconditioned Random $3$-Regular Graph Maximum-Cut Problems
Number of QAOA layers Number of nonzero terms for $\boldsymbol{p\ll\ln N}$ Number of terms to precondition for $\boldsymbol{p\ll\ln N}$
Original 1.5N
$p=1$ $4.500008(8)N$ $11.2(4)$
$p=2$ $22.5000(2)N$ $740(17)$
$p=3$ $94.500(2)N$ $18,581(231)$
$p=4$ $382.45(3)N$ $360,937(2,832)$
@@ Appendix E Terms in Quantum-Preconditioned Random $3$-Regular Graph Maximum-Cut Problems
Table 2: Asymptotic behaviors from Fig. 12. Numbers of nonzero terms for $p\ll\ln N$ extracted by a linear fit of the four largest problem sizes. Numbers of terms to precondition for for $p\ll\ln N$ correspond to the average value of the largest problem size with $N=2^{16}$ variables.
@@
Random $3$-regular graph maximum-cut problems are sparse: The upper triangle of the adjacency matrix $\mathsf{W}$ contains $23N/$ 2 entries. The QAOA with $p$ layers induces nonzero correlations $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle$ at a distance $2p$ between variables $i$ and $j$ via a light cone. This results in additional nonzero terms in the preconditioned problem when compared to the original one. We show in Fig. 12a the number of terms $n$ as a function of the number of variables $N$ for various values of $p$. We observe two asymptotic behaviors: $n\sim O(N)$ for $p\ll\ln N$ and $n\sim O(N^{2})$ for $p\gg\ln N$, understood through the topology of random $3$-regular graphs having an average distance between two variables going as $\ln N$.
@@
Although the number of nonzero terms grows with $N$, only an $N$-independent number $O(\exp p)$ of terms needs to be individually preconditioned for $p\ll\ln N$. This is because asymptotically with $N$, most of the correlations $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle$ are identical: They correspond to light-cone-induced tree subgraphs. A tree is straightforwardly detected as a graph for which the number of edges $n$ is $N=n+1$ and its corresponding correlation value $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle$ needs only be computed once and cached.
@@
We report in Tab. 2 asymptotic numerical values. We conjecture that the prefactor in the $O(N)$ scaling for the average number of nonzero terms relates to a sequence of integers $\{3,~{}9,~{}45,~{}189,~{}765,~{}\ldots\}$, which is, however, not referenced on the online encyclopedia of integer sequences (OEIS).
@@
Appendix F Semi-Analytical Expression for the Correlation Matrix at $p=1$
The proposed quantum preconditioning can be performed at $p=1$ for arbitrary $N$-variable problems classically and semi-analytically using a back-propagation technique for evaluating expectation values $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle_{p=1}$ as a function of the QAOA parameters $\gamma_{1}\equiv\gamma$ and $\beta_{1}\equiv\beta$, as defined in Eq. (3). For an arbitrary objective function (arbitrary adjacency matrix $\mathsf{W}$) of the form of Eq. (1), the complexity for computing the full correlation matrix is $O(N^{3})$. We simply state the result derived in Ref. [49] for completeness:
@@ Appendix F Semi-Analytical Expression for the Correlation Matrix at $p=1$
×$\times\Biggl{[}\prod\nolimits_{k\neq i,j}\cos\bigl{(}\gamma\mathsf{W}_{ik}\bigr{)}+\prod\nolimits_{k\neq i,j}\cos\bigl{(}\gamma\mathsf{W}_{jk}\bigr{)}\Biggr{]}$
@@
We use Eq. (19) for preconditioning the Sherrington-Kirkpatrick spin glass problems studied in the main text.
@@
Appendix G Performance Versus Number of Iterations
G.1 Random $3$-Regular Graph Maximum-Cut Problems
We investigate the performance for the maximum-cut problem on $N$-variable random $3$-regular graphs based on the original and quantum-preconditioned problems. We consider SA and the BM solvers on problem sizes ranging from $N=32$ to $N=4,096$. For each problem size, we average the performance over $200$ randomly generated instances. The preconditioning is performed using the light-cone technique described in Sec. III.2.1.
We report data in Figs. 13 and 14. We observe an advantage for the preconditioned problems against the original one as a function of the number of iterations. The advantage is enhanced by going from $p=1$ to $p=2$. We observe that for a large number of iterations, the approximation ratio for the quantum preconditioned problem at $p=2$ saturates. This is because the original and preconditioned problems do not share the same optimal solutions. The approximation ratio saturates to an $N$-dependent value compatible with $1-\alpha_{\textrm{saturation}}\sim N^{-1}$ up to $N=1,024$ (not enough iterations were performed to observe the saturation for larger values of $N$).
The number of iterations can be converted to a run-time following Appendix H. We presented run-time data in the main text.
@@ G.1 Random $3$-Regular Graph Maximum-Cut Problems
Figure 15: Average performance (approximation ratio versus the number of iterations) for $N$-variable Sherrington-Kirkpatrick spin glass problems via the classical simulated annealing (SA) and Burer-Monteiro (BM) solvers based on the original and quantum-preconditioned $(p=1)$ problems. (A) $N=32$ variables. (B) $N=64$ variables. (C) $N=128$ variables. (D) $N=256$ variables. (E) $N=512$ variables. (F) $N=1,024$ variables. Each data point is averaged over $200$ randomly generated problem instances. Error bars indicate the standard error of the mean.
@@
Figure 16: Average run-time per term in seconds (based on 64Gb MacBook Pro with an Apple M1 Max chip) versus the number of iterations for original and preconditioned problems. (A) Simulated annealing (SA) solver for random $3$-regular graph maximum-cut problems. (B) Burer-Monteiro (BM) solver for random $3$-regular graph maximum-cut problems. (C) SA solver for Sherrington-Kirkpatrick spin glasses. (D) BM solver for Sherrington-Kirkpatrick spin glasses. Error bars indicate the standard error of the mean. Bold lines are fits according to the main text.
G.2 Sherrington-Kirkpatrick Spin Glasses
@@
Next, we provide additional data on Sherrington-Kirkpatrick spin glasses. We investigate the performance of SA and the BM solver on the original and quantum preconditioned ($p=1$) problems. Problem sizes ranging from $N=32$ to $N=1,024$ are considered ($N=2,408$ was considered in the main text in Fig. 3). We report data in Fig. 15. We find that an advantage for the preconditioned problem for the BM solver only appears for large problem sizes. We highlight this advantage window in yellow. The average approximation ratio saturation value observed for a large number of iterations, $\alpha\simeq 99.3\%$, seems to be roughly independent of $N$.
Appendix H From Number of Iterations to Run-time
@@
Instead of run-time, the performance of the SA and BM solvers can be estimated through the number of iterations, which is a platform-independent number.
@@
The run-time $t$ of a solver follows asymptotically with $N$
@@
where $n$ is the number of terms in the problem, $n_{\textrm{iter}}$ the number of iterations. $A$ and $B$ are solver-, platform-, and problem-dependent constants. In the case of SA, the constant $B$ is nonzero and corresponds to a one-time $O(n)$ heuristic for determining the temperature schedule based on the problem’s terms (see Appendix D). It is zero for the BM solver, which doesn’t have any hyperparameters.
We show the relation between run-time and number of iterations in Fig. 16 for the random $3$-regular graph maximum-cut problems and Sherrington-Kirkpatrick spin glasses for SA and the BM solvers on the original and preconditioned problems. The run-time is based on a 64Gb MacBook Pro with an Apple M1 Max chip. In the case of $N$-variable $3$-regular graphs, we have $n=1.5N$ terms in the original problem, $n\simeq 4.5N$ at $p=1$, and $n\simeq 22.5N$ at $p=2$ in the large $N$ limit (see Appendix E). For Sherrington-Kirkpatrick spin glasses, the number of terms is $n=N(N-1)/2$ in all cases. We reported the average approximation ratio as a function of both the number of iterations and run-time for the grid energy problem in Fig. 4.

### v2

@@ E.2 Burer-Monteiro solver
The BM solver eliminates the positive semidefinite constraint of the GW algorithm for the maximum-cut problem. The GW algorithm is based on semidefinite programming methods and the solver guarantees the highest possible approximation ratio $\alpha\simeq 0.878$ [10, 11]. Thus, the BM solver involves a nonconvex objective function, which may result in multiple local nonglobal minima. In practice, this is not an issue, and the BM solver is considered one of the best heuristics to date for the maximum-cut problem [12]. Its foundations are still intensively studied by the community; see, e.g., [167, 168] and references therein.
@@
We employ the MQLib implementation of the BM solver [12], written in C++ for efficiency. This implementation does not readily expose any hyperparameters for the solver and we keep the default values. This is why, unlike for SA, the run-time is strictly proportional to the number of iterations (see Appendix K). Our only modification of the MQLib code exposes the number of iterations performed by the solver in addition to its run-time.
Appendix F Quantum Preconditioning Performance Away from Near-Optimal QAOA Angles
We investigate the performance of the proposed quantum-preconditioning approach away from near-optimal QAOA angles. We perform numerical simulations on $N=64$ SK problem instances at $p=1$ with QAOA angles defined as
$\gamma=\bigl(\varepsilon\cos\theta+0.5\bigr)/\sqrt{N},~~\beta=\varepsilon\sin\theta+\pi/8,$ (20)
where $\varepsilon$ measures the Euclidean distance to the near-optimal tabulated QAOA angles and $\theta$ is a random number uniformly drawn from the range $0,2\pi]$. For $\varepsilon=0$, one recovers the near-optimal tabulated angles [27].
In Fig. 12, we show the effect of $\varepsilon>0$ relative to the $\varepsilon=0$ case on the performance of simulated annealing versus the number of iterations, as measured by $\alpha(\varepsilon)/\alpha(\varepsilon=0)$, where $\alpha$ is the average approximation ratio over problem instances and samples. We observe that the quantum-preconditioning approach is quite robust to variations of QAOA angles. The sensitivity to $\varepsilon$ for larger $p$ remains open.
Finally, we note that for random $3$-regular graphs, the tabulated angles that we used throughout this work are not optimal on an instance-by-instance basis but are only near-optimal on average [51], meaning that they correspond to an effective finite $\varepsilon$ value; however, this near-optimality of the angles was enough to provide a quantum-inspired advantage via quantum preconditioning.
@@ Appendix F Quantum Preconditioning Performance Away from Near-Optimal QAOA Angles
Appendix G Terms in Quantum-Preconditioned Random $3$-Regular Graph Maximum-Cut Problems
Figure 13: (A) Average number of nonzero terms in quantum-preconditioned random $3$-regular graph maximum-cut problems for various values of $p$ as a function of the number of variables NN. (B) Average number of terms to precondition in quantum-preconditioned random $3$-regular graph maximum-cut problems for various values of $p$ as a function of the number of variables NN. It is smaller than or equal to the number of nonzero term via caching the preconditioned value of light-cone-induced tree subgraphs. Each data point is averaged over $200$ randomly generated problem instances. Error bars indicate the standard error of the mean.
Number of
QAOA layers
Number of
nonzero terms
for $\boldsymbol{p\ll\ln N}$
Number of
terms to precondition
for $\boldsymbol{p\ll\ln N}$
Original
1.5N
$p=1$
$4.500008(8)N$
$11.2(4)$
$p=2$
$22.5000(2)N$
$740(17)$
$p=3$
$94.500(2)N$
$18,581(231)$
$p=4$
$382.45(3)N$
$360,937(2,832)$
@@ Appendix G Terms in Quantum-Preconditioned Random $3$-Regular Graph Maximum-Cut Problems
Table 2: Asymptotic behaviors from Fig. 13. Numbers of nonzero terms for $p\ll\ln N$ extracted by a linear fit of the four largest problem sizes. The numbers of terms to precondition for $p\ll\ln N$ correspond to the average value of the largest problem size with $N=2^{16}$ variables.
@@
Random $3$-regular graph maximum-cut problems are sparse: the upper triangle of the adjacency matrix $\mathsf{W}$ contains $3N/2$ entries. The QAOA with $p$ layers induces nonzero correlations $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle$ at a distance $2p$ between variables $i$ and $j$ via a light cone. This results in additional nonzero terms in the preconditioned problem when compared to the original one. In Fig. 13a, we show the number of terms $n$ as a function of the number of variables $N$ for various values of $p$. We observe two asymptotic behaviors: $n\sim O(N)$ for $p\ll\ln N$ and $n\sim O(N^{2})$ for $p\gg\ln N$, understood through the topology of random $3$-regular graphs having an average distance between two variables going as $\ln N$.
@@
Although the number of nonzero terms grows with $N$, only an $N$-independent number $O(\exp p)$ of terms needs to be individually preconditioned for $p\ll\ln N$. This is because asymptotically with $N$, most of the correlations $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle$ are identical: they correspond to light-cone-induced tree subgraphs. A tree is straightforwardly detected as a graph for which the number of edges $n$ is $N=n+1$ and its corresponding correlation value $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle$ only needs to be computed once and cached.
@@
In Tab. 2, we report asymptotic numerical values. We conjecture that the prefactor in the $O(N)$ scaling for the average number of nonzero terms relates to a sequence of integers $\{3,~9,~45,~189,~765,~\ldots\}$, which is, however, not referenced in the online encyclopedia of integer sequences (OEIS).
@@
Appendix H Semi-Analytical Expression for the Correlation Matrix at $p=1$
The proposed quantum preconditioning can be performed at $p=1$ for arbitrary $N$-variable problems classically and semianalytically using a back-propagation technique for evaluating expectation values $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle_{p=1}$ as a function of the QAOA parameters $\gamma_{1}\equiv\gamma$ and $\beta_{1}\equiv\beta$, as defined in Eq. (3). For an arbitrary objective function (arbitrary adjacency matrix $\mathsf{W}$) of the form of Eq. (1), the complexity for computing the full correlation matrix is $O(N^{3})$. We simply state the result derived in Ref. [48] for completeness:
@@ Appendix H Semi-Analytical Expression for the Correlation Matrix at $p=1$
$\times\Biggl[\prod\nolimits_{k\neq i,j}\cos\bigl(\gamma\mathsf{W}_{ik}\bigr)+\prod\nolimits_{k\neq i,j}\cos\bigl(\gamma\mathsf{W}_{jk}\bigr)\Biggr]$
@@
We use Eq. (21) for preconditioning the Sherrington-Kirkpatrick spin-glass problems studied in the main text.
@@
Appendix I Performance Against Approximate Classical Tensor Network Emulations
@@
Figure 16: Performance of simulated annealing for quantum-preconditioned SK problems with approximate MPS emulations. Ratio $1-C_{\chi}/C_{\chi\to\infty}$ against $\ln\chi/N$, where $\chi$ is the MPS bond dimension and $N$ is the number of variables. The value $C_{\chi}$ is the average objective value of $10^{3}$ solutions sampled with simulated annealing. Each data point is averaged over $50$ randomly generated problem instances. Error bars indicate the standard error of the mean.
To complement the discussion of Sec. IV.2.3, we perform classical matrix-product-state (MPS) emulations of the $p=1$ QAOA used to precondition an ensemble of $N$-variable SK problems. We employ the Python library quimb [169] for the MPS emulations. The QAOA is emulated approximately, with the emulation quality controlled by the MPS bond dimension $\chi$. Then, we use simulated annealing with a fixed number of $10^{4}$ iterations to evaluate the performance of the classical solver against approximately quantum-preconditioned problems.
Inspired by Ref. [73], and for each problem instance, we consider the average objective value $C_{\chi}$ over $10^{3}$ solutions sampled by simulated annealing. The comparison point is with respect to the solutions returned from perfectly quantum-preconditioned problems, i.e., $C_{\chi\to\infty}$, which we can efficiently access at $p=1$ based on Sec. H.
In Fig. 16, we observe a data collapse for $C_{\chi}/C_{\chi\to\infty}$ as a function of the rescaled variable $\ln\chi/N$. This behavior for the output of simulated annealing with quantum preconditioning was also observed for the standard QAOA [73], on which the quantum-preconditioning step is based.
For example, one needs $\ln\chi/N\gtrsim 0.1$ for the average quality solutions of approximately quantum-preconditioned simulated annealing to be $90\%$ as good as the exactly ($\chi\to+\infty$) preconditioned one.
Appendix J Performance Versus Number of Iterations
J.1 Random $3$-regular graph maximum-cut problems
We investigate the performance for the maximum-cut problem on $N$-variable random $3$-regular graphs based on the original and quantum-preconditioned problems. We consider SA and the BM solver on problem sizes ranging from $N=32$ to $N=4,096$. For each problem size, we average the performance over $200$ randomly generated instances. The preconditioning is performed using the light-cone technique described in Sec. III.1.2.
We report data in Figs. 14 and 15. We observe an advantage for the preconditioned problems against the original one as a function of the number of iterations. This advantage is enhanced by going from $p=1$ to $p=2$. We observe that for a large number of iterations, the approximation ratio for the quantum-preconditioned problem at $p=2$ saturates. This is because the original and preconditioned problems do not share the same optimal solutions. The approximation ratio saturates to an $N$-dependent value compatible with $1-\alpha_{\textrm{saturation}}\sim N^{-1}$ up to $N=1,024$ (not enough iterations were performed to observe the saturation for larger values of $N$).
The number of iterations can be converted to a run-time following Appendix K. We present run-time data in the main text.
Figure 17: Average performance (approximation ratio versus the number of iterations) for $N$-variable Sherrington-Kirkpatrick spin glass problems via the classical SA and BM solvers based on the original and quantum-preconditioned $(p=1)$ problems. (A) $N=32$ variables. (B) $N=64$ variables. (C) $N=128$ variables. (D) $N=256$ variables. (E) $N=512$ variables. (F) $N=1,024$ variables. Each data point is averaged over $200$ randomly generated problem instances. Error bars indicate the standard error of the mean.
@@
Figure 18: Average run-time per term in seconds (based on 64GB MacBook Pro with an Apple M1 Max chip) versus the number of iterations for original and preconditioned problems. (A) SA solver for random $3$-regular graph maximum-cut problems. (B) BM solver for random $3$-regular graph maximum-cut problems. (C) SA solver for Sherrington-Kirkpatrick spin glasses. (D) BM solver for Sherrington-Kirkpatrick spin glasses. Error bars indicate the standard error of the mean. Bold lines are fits according to the main text.
J.2 Sherrington-Kirkpatrick spin glasses
@@
Next, we provide additional data on Sherrington-Kirkpatrick spin glasses. We investigate the performance of SA and the BM solver on the original and quantum-preconditioned ($p=1$) problems. Problem sizes ranging from $N=32$ to $N=1,024$ are considered ($N=2,408$ was considered in the main text in Fig. 3). We report data in Fig. 17. We find that an advantage for the preconditioned problem for the BM solver only appears for large problem sizes. We highlight this advantage window in yellow. The average approximation ratio saturation value observed for a large number of iterations, $\alpha\simeq 99.3\%$, seems to be roughly independent of $N$.
Appendix K From Number of Iterations to Run-time
@@
Instead of using the run-time, the performance of the SA and BM solvers can be estimated through the number of iterations, which is a platform-independent number.
@@
The run-time $t$ of a solver scales asymptotically with $N$ as
@@
where $n$ is the number of terms in the problem and $n_{\textrm{iter}}$ is the number of iterations. Here, $A$ and $B$ are solver-, platform-, and problem-dependent constants. In the case of SA, the constant $B$ is nonzero and corresponds to a one-time $O(n)$ heuristic for determining the temperature schedule based on the problem’s terms (see Appendix E). This is zero for the BM solver, which does not have any hyperparameters.
We show the relation between the run-time and number of iterations in Fig. 18 for the random $3$-regular graph maximum-cut problems and Sherrington-Kirkpatrick spin glasses for SA and the BM solver on the original and preconditioned problems. The run-time is based on a 64GB MacBook Pro with an Apple M1 Max chip. In the case of $N$-variable 33-regular graphs, we have $n=1.5N$ terms in the original problem, $n\simeq 4.5N$ at $p=1$, and $n\simeq 22.5N$ at $p=2$ in the large-$N$ limit (see Appendix G). For Sherrington-Kirkpatrick spin glasses, the number of terms is n=N(N1)/2n=N(N-1)/2 in all cases. We reported the average approximation ratio as a function of both the number of iterations and run-time for the grid energy problem in Fig. 4.